OpenStudy (anonymous):

Solve this quadratic equation using the quadratic formula. 2x2 - 2x = 1

OpenStudy (luigi0210):

Do you know the quadratic formula?

OpenStudy (anonymous):

yes

pooja195 (pooja195):

\[\huge~\rm~~x=\frac{ -b~\pm \sqrt{b^2-4ac} }{ 2a } \]

pooja195 (pooja195):

lst your abc values

OpenStudy (luigi0210):

Okay, so in order to solve it using the formula you need your function in \(ax^2+bx+c =0\) So set your function equal to 0 first.